In a class of 70 students 6 offer Economics only, 18 offer Geography and economics but not Mathematics, 36 offer Economics and Mathematics, 53 offer Economics, 50 offer Geography and 34 offer Mathematics. All students offer at least one subject.
(a)Determine the number of students who offer Maths only
(b) Determine the number of students who Geography only

Answer by math_tutor2020(3817)


Venn Diagram

There are 8 distinct regions
a = Econ only
b = Econ and Geo, but not Math
c = Geo only
d = Econ and Math, but not Geo
e = all 3
f = Geo and Math, but not Econ
g = Math only
h = neither of the 3 classes mentioned

Given facts
NumberFact
170 students
26 offer Economics only
318 offer Geography and economics but not Mathematics
436 offer Economics and Mathematics
553 offer Economics
650 offer Geography
734 offer Mathematics.
8All students offer at least one subject.


Fact 2 leads directly to a = 6
Fact 3 tells us b = 18

fact4 says that d+e = 36

we have
a = 6
b = 18
d+e = 36
so,
a+b+d+e = 6+18+36 = 60

But this sum is larger than the number of people taking Econ (see fact 5). We run into a contradiction.

Put another way:
fact5 says that a+b+d+e = 53
Let's plug a = 6 and b = 18 into that and isolate d+e
a+b+d+e = 53
6+18+d+e = 53
24+d+e = 53
d+e = 53-24
d+e = 29
This clashes with fact 4 that d+e = 36

Either way we run into a contradiction.
It's likely your teacher made a typo somewhere when coming up with the numbers to this problem.
I would ask him/her for further clarification.